OpenStudy (anonymous):

I am studying for a test tomorrow and i would greatly appreciate it if someone would help me find the y intercept of 3y=15x-9

jimthompson5910 (jim_thompson5910):

To find the y-intercept, you plug in x = 0 and solve for y 3y=15x-9 3y=15(0)-9 ... replace x with 0 3y = 0 - 9 3y = -9 y = ???

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

what did you get for y?

OpenStudy (anonymous):

Well, if i read your equation right it is -9

jimthompson5910 (jim_thompson5910):

3y = -9 but we want to know what y is (not 3y)

jimthompson5910 (jim_thompson5910):

if 3 times y = -9, then y = ???

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

So the. What would 2x+8y=12

OpenStudy (anonymous):

Then*

jimthompson5910 (jim_thompson5910):

good, the y-intercept is -3 which means the graph crosses the y axis at the point (0,-3)

jimthompson5910 (jim_thompson5910):

to find the y-intercept of 2x+8y=12, you plug in x = 0 and solve for y 2x+8y=12 2(0)+8y=12 ... ... ... y = ???

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

that's the x-intercept actually

jimthompson5910 (jim_thompson5910):

because plugging in y = 0 gives you 2x+8y=12 2x+8(0)=12 2x = 12 x = 12/2 x = 6

OpenStudy (anonymous):

I am sosorry, you are right, i looked badly

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

So then the y intercept would be 8

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

if 8y = 12, then y = ???

jimthompson5910 (jim_thompson5910):

hopefully you see how I'm getting 8y = 12

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

I plugged in x = 0 into 2x+8y=12

jimthompson5910 (jim_thompson5910):

2x+8y=12 2(0)+8y=12 0 + 8y = 12 8y = 12

OpenStudy (anonymous):

Yes, i get that already, but the thing is how can you grab 8 times y equals 12

jimthompson5910 (jim_thompson5910):

8y means "8 times y"

jimthompson5910 (jim_thompson5910):

what undoes multiplication?

OpenStudy (anonymous):

So the oppposite of multiplication

OpenStudy (anonymous):

Or just minus

jimthompson5910 (jim_thompson5910):

no, division actually

jimthompson5910 (jim_thompson5910):

you divide both sides by 8 to completely isolate y

OpenStudy (anonymous):

Because the oppposite is divsion

jimthompson5910 (jim_thompson5910):

yes, division undoes multiplication (by definition)

OpenStudy (anonymous):

So we end up with?

jimthompson5910 (jim_thompson5910):

you tell me

jimthompson5910 (jim_thompson5910):

8y = 12 divide both sides by 8

OpenStudy (anonymous):

How do i do that, 8(8y=12)

OpenStudy (anonymous):

Like that?

jimthompson5910 (jim_thompson5910):

8y/8 turns into 1y/1 or just y on the left side

jimthompson5910 (jim_thompson5910):

12/8 = ???

OpenStudy (anonymous):

1.5

jimthompson5910 (jim_thompson5910):

or 3/2

jimthompson5910 (jim_thompson5910):

the y-intercept is 3/2 or 1.5

OpenStudy (anonymous):

Does the slash mean division for you people

jimthompson5910 (jim_thompson5910):

"you people"?

jimthompson5910 (jim_thompson5910):

sounds a bit rude

OpenStudy (anonymous):

You guys

jimthompson5910 (jim_thompson5910):

but yeah, / means division

jimthompson5910 (jim_thompson5910):

2/3 means "2 divided by 3" and it's also fraction notation

OpenStudy (anonymous):

Are you a teacher by any chance

jimthompson5910 (jim_thompson5910):

not yet, but I'm on my way to becoming one

OpenStudy (anonymous):

Thats great, keep on it

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

Bye

jimthompson5910 (jim_thompson5910):

bye